# Towards Understanding Link Predictor Generalizability Under Distribution Shifts

### v2

@@ Abstract.
State-of-the-art link prediction (LP) models demonstrate impressive benchmark results. However, popular benchmark datasets often assume that training, validation, and testing samples are representative of the overall dataset distribution. In real-world situations, this assumption is often incorrect; uncontrolled factors lead new dataset samples to come from a different distribution than training samples. Additionally, the majority of recent work with graph dataset shift focuses on node- and graph-level tasks, largely ignoring link-level tasks. To bridge this gap, we introduce a novel splitting strategy, known as LPShift, which utilizes structural properties to induce a controlled distribution shift. We verify LPShift’s effect through empirical evaluation of SOTA LP models on 16 LPShift variants of original dataset splits, with results indicating drastic changes to model performance. Additional experiments demonstrate graph structure has a strong influence on the success of current generalization methods. Source Code Available Here: LPShift
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Ranking
@@ 1. Introduction
Furthermore, while numerous methods work to account for distribution shifts within graph machine learning (Li et al., 2022b), there remains little work doing so for LP. Specifically, we observe that (1) No LP Benchmark Datasets: Current graph benchmark datasets designed with a quantifiable distribution shift are focused solely on the node and graph tasks (Zhou et al., 2022a; Ding et al., 2021), with no datasets available for LP. (2) Absence of Foundational Work: There is limited existing work for distribution shifts relevant to LP (Zhang et al., 2022). Current methods are primarily focused on detecting and alleviating anomalies within node- and graph-level tasks (Jin et al., 2022; Bevilacqua et al., 2021; Gao et al., 2023; Wu et al., 2024; Guo et al., 2023; Wu et al., 2023; Sui et al., 2024; Li et al., 2022b). Additionally, few methods are designed for aiding LP generalization in any setting (Dong et al., 2022; Zhao et al., 2022; Zhang et al., 2022; Zhou et al., 2022a). Also, other LP generalization methods which are theorized to improve performance in shifted scenarios remain crucially untested (Singh et al., 2021; Wang et al., 2023b).
To tackle these problems, this work proposes the following contributions:
  • Creating Datasets with Meaningful Distribution Shifts. LP requires pairwise structural considerations (Liben-Nowell and Kleinberg, 2003; Mao et al., 2024a). Additionally, when considering realistic settings (Li et al., 2024) or distribution shift (Zhu et al., 2024), GNN4LP models perform poorly relative to models used in graph (Wei et al., 2021; Yuan et al., 2021) and node classification (Shi et al., 2023; Zhao et al., 2023). To better understand distribution shifts, we use key structural LP heuristics to split the links into train/validation/test splits via LPShift. By applying LPShift to generate dataset splits, we induce shifts in the underlying distribution of the links, thereby affecting formation (Mao et al., 2024a). Theoretical justification is provided in Section 4.
@@
  • Benchmarking Current LP Methods. GNN4LP models demonstrate a strong sensitivity to LPShift when generalizing to the synthetic dataset splits. Despite the existence of LP generalization methods, such as FakeEdge (Dong et al., 2022) and Edge Proposal Sets (Singh et al., 2021), there remains little work benchmarking link-prediction models under distribution shifts (Ding et al., 2021; Dong et al., 2022; Zhu et al., 2024). This lack of benchmarking contributes to a gap in understanding, impeding the capabilities of LP models to generalize. This work quantifies the performance of SOTA LP models under 16 unique LPShift scenarios and provides analysis as a foundation for improving LP model generalization. We further quantify the effects of LP and graph-specific generalization methods, finding that they also struggle to generalize with differing structural shifts.
@@
The remainder of this paper is structured as follows. In Section 2, we provide background on the heuristics, models, and generalization methods used in LP. In Section , we detail how the heuristics, theoretical perspectives, and HeaRT (Li et al., 2024) relate to our proposed splitting strategy and formally introduce LPShift. Lastly, in Section 5, we benchmark a selection of LP models and generalization methods on LPShift, followed by analysis to understand the effects of this new strategy.
@@ 3. Benchmark Dataset Construction
7:     $\Psi(u,v)=h(u,v)$triangleright Score edge with neighborhood heuristic
8:     if $h(u,v)\leq i_{train}$ then $\triangleright$ Train Split
@@
10:     else if $h(u,v)>i_{train}\text{ and }h(u,v)\leq i_{valid}$ then $\triangleright$ Valid split
@@
3.1. Types of Distribution Shifts
We induce distribution shifts by splitting the links based on key structural properties affecting link formation and thereby LP. We consider three type of metrics: Local structural information, Global structural information, and Preferential Attachment. Recent work by (Mao et al., 2024b) has shown the importance of local and global structural information for LP. Furthermore, due to the scale-free nature of many real-world graphs relates to link formation (Barabási and Albert, 1999), we also consider Preferential Attachment. A representative metric is then chosen for each of the three types, shown as follows:
[…]
3.2. Dataset Splitting Strategy
In the last subsection we described the different types of metrics to induce distribution shifts for LP. The metrics cover fundamental structural properties that influence the formation of new links. We now describe how we use these measures to split the dataset into train/validation/test splits to induce such shifts.
In order to build datasets with structural shift, we apply a given neighborhood heuristic to score each link. This score is then compared to a threshold ($i_{train},i_{valid})$ to categorize a link as a different sample. As denoted in Alg. 1, the heuristic score of the link $(u,v)$ is $h(u,v)$. The link falls into: training when $h(u,v)<i_{train}$, validation when $i_{train}<h(u,v)\leq i_{valid}$, and testing when $h(u,v)>i_{valid}$. The new training graph is constructed from the original OGB dataset (Hu et al., 2020). Validation and testing samples are removed from the new training graph to prevent test-leakage and limited to 100-thousand edges maximum. The full algorithm is detailed in Algorithm 1 with additional details in Appendix C.
With Figure 2, we provide a small example of how splits are produced by our proposed splitting strategy. Specifically, Figure 2(a) demonstrates an outcome of the CN split labelled “CN - 1,2” where sampled edges pulled from the: black-dotted line = training (no CNs), red-dotted line = validation, (1 CN), and blue-dotted line = testing ($\geq$2 CNs). See Appendix C for information on Figure 2(b) and Figure 2(c).
(a) CN
(b) PA
(c) SP
[…]
To test how different LPShift thresholds impact performance, we adjust the $i_{train}$ and $i_{valid}$ thresholds to produce 3 varied CN and PA splits; as well as 2 varied SP splits. The variations in split thresholds were chosen based on two conditions. 1) Structural information within splits varies due to user-defined thresholds. In the ”Forward” scenario (i.e. CN - 1,2), splits are given increasingly more structural information between training to validation, giving the model an easier time generalizing under testing. For the ”Backward” scenario (i.e, CN - 2,1), stricter thresholds mean less structural information between validation to testing; making generalization difficult. 2) The final dataset split contains a sufficient number of samples. Each LPShift split requires enough split samples to allow model generalization. Given the limited number of SP split samples, the SP splits were limited to 2 variants. See Appendix C for more algorithmic details, and Appendix E for more information on LPShift’s efficiency and usefulness.
@@ 3.3. Why use HeaRT with LPShift?
It is common practice for most link prediction baselines to randomly-sample all negatives samples. However, random sampling is shown to be unrealistic, given that it is trivially-easy for a GNN4LP model trained via supervised learning to distinguish positive and negative samples based solely on the sample’s structure (Li et al., 2024). We demonstrate an example of this in the right subplot of Figure 1, where negative samples overwhelmingly contain 0 Common Neighbors versus HeaRT, whose negative samples are more closely-aligned with their respective positive samples. LPShift’s ability to induce distribution shift is explored further in Appendix Section G
@@ 4. Why LPShift? A Theoretical Perspective
As neural models with backbone GNNs, GNN4LP relies on assumptions about the dataset’s underlying data distributions to learn representations for effective link prediction. Due to these assumptions, GNN4LP capture structural heuristics associated with link formation, such as LPFormer estimating Katz-Index (Shomer et al., 2024). LPShift undermines these assumptions through direct manipulation of the dataset sample distribution. We provide a theoretical perspective on this effect below:
@@
The supervised nature of training neural link prediction models with GNN backbones means that message-passing across training samples allows GNN4LP to observe and then model link formation to enhance downstream performance (Bronstein et al., 2021; Zečević et al., 2021). However, Theorem 4.2 functions as a way of conditioning link formation dynamics within a given graph dataset so that necessary subsets are obscured. In effect, meaning the model gets a limited view on pairwise information contained within the dataset. As such, LPShift tests the hypothesis: Can GNN4LP model structural heuristics for effective downstream performance in scenarios with limited structural information? We verify this effect empirically in Section 5.
@@ 5. Experiments
To bridge the gap for GNN4LP generalizing under distribution shifts, this work addresses the following questions: (RQ1) Is the distribution shift induced by LPShift significant? (RQ2) Can SOTA GNN4LPs generalize under our proposed distribution shifts? (RQ3) Can current generalization methods further boost the performance of current methods? (RQ4) What components of the proposed distribution shift are affecting the LP model’s performance?
@@ 5.1. Experimental Setup
Datasets: We consider 16 “Forward” and “Backward” LPShift splits for the following 3 OGB datasets (Hu et al., 2020): ppa, ddi, and collab, for a total of 48 tested splits. The resulting datasets represent tasks in three separate domains and three shifted scenarios, allowing a comprehensive study of LP generalization under distribution shift. For all datasets, we create multiple splits corresponding to each structural property detailed in Section 3.2. For the “Forward” split, denoted as $(X,Y,Z)$, an increase in $Y$ and $Z$ indicates more structural information available to the training adjacency matrix. The “Backward” split swaps the training and testing splits from their counterpart in the “Forward” split, resulting in the training adjacency matrix losing access to structural information as $X$ and $Y$ increase. See Appendix J for more details.
@@
Graph-Specific Generalization Methods: We also test the performance of BUDDY with multiple generalization techniques. This includes DropEdge (Rong et al., 2020), which randomly removes a portion of edges from the training adjacency matrix. Edge Proposal Sets (EPS) (Singh et al., 2021), which uses paired LP models to filter and rank top-k edges to insert as Common Neighbors in the training adjacency matrix. Lastly, we consider Topological Concentration (TC) (Wang et al., 2023b), which re-weights the edges within the training adjacency matrix based on the structural information.
Traditional Generalization Methods: The final round of tests applies: IRM (Arjovsky et al., 2019), VREx (Krueger et al., 2021), GroupDRO (Sagawa et al., 2019), DANN (Ganin et al., 2016), and DeepCORAL (Sun and Saenko, 2016) as generalization methods to GCN. Given that LPShift splits graph data directly on the sample structure, all tests with traditional generalization methods treat training samples as an environment, =$\mathcal{E}=\{e_{1},...,e_{N}\}$ composed of $N$ environmental subsets, where each subset is binned based on their neighborhood heuristic; as demonstrated for CNs in Figure 3. This binning of environmental subsets forces traditional generalization methods to target graph structure, controlling for the method’s ability to improve performance under structural shift induced by LPShift.
Figure 3. The original environment of dataset samples split into subsets by their CN heuristic scores.
@@
The original environment of dataset samples split into subsets by their CN heuristics.
@@
Hyperparameters: All methods were tuned on permutations of learning rates in $\{1e^{-2},1e^{-3}\}$ and dropout in $\{0.1,0.3\}$. Each model was trained and tested over five seeds to obtain the mean and standard deviations of their results. Given the significant time complexity of training and testing on the customized ogbl-ppa datasets, NCNC and LPFormer were tuned on a single seed, followed by an evaluation of the tuned model on five separate seeds. Additional hyperparameter tuning details are included in Appendix Section J.
5.2. Results for GNN4LP
In order to provide a unified perspective on how distribution shift affects link prediction models, each GNN4LP method was trained and tested across five seeded runs on versions of ogbl-collab, ppa, and ddi split by: Common Neighbors, Shortest-Path, and Preferential-Attachment. Examining the results, we have the following three key observations.
Observation 1: Drastic Changes in Model Performance. As shown in Table 1, the majority of synthetic splits generated by LPShift reduce model performance significantly below the original HeaRT standard (Li et al., 2024). This is especially pronounced in denser graphs like ogbl-ppa and ogbl-ddi, indicating that LPShift affects the varied structures of graph data which then reduces downstream perfomance. The only exception is ogbl-collab, where: RA, NCNC, LPFormer, and SEAL experience measurable improvements over the HeaRT baseline. Given ogbl-collab’s lack of density, relative to ogbl-ppa and ogbl-ddi, there are considerations about how LPShift induces a covariate shift to features (Koh et al., 2021), especially where structure is correlated to the feature distribution.
Additional analysis shown in Figure 6 and 7 indicates how dataset structure relates to the performance effects of LPShift. Furthermore, to quantify the extent of how the node features correlate with LPShift, we measure the cosine similarity of each split’s feature distribution. These results are shown in Appendix D.

### v3

@@ Abstract.
State-of-the-art link prediction (LP) models demonstrate impressive benchmark results. However, popular benchmark datasets often assume that training, validation, and testing samples are representative of the overall dataset distribution. In real-world situations, this assumption is often incorrect; uncontrolled factors lead new dataset samples to come from a different distribution than training samples. Additionally, the majority of recent work with graph dataset shift focuses on node- and graph-level tasks, largely ignoring link-level tasks. To bridge this gap, we introduce a novel splitting strategy, known as LPShift, which utilizes structural properties to induce a controlled distribution shift. We verify LPShift’s effect through empirical evaluation of SOTA LP models on 16 LPShift variants of original dataset splits, with results indicating drastic changes to model performance. Additional experiments demonstrate graph structure has a strong influence on the success of current generalization methods. The source code is available here https://github.com/revolins/LPShift.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3737434isbn: 979-8-4007-1454-2/2025/08ccs: Mathematics of computing Graph algorithmsccs: Computing methodologies Supervised learning
@@ 1. Introduction
Furthermore, while numerous methods work to account for distribution shifts within graph machine learning (Li et al., 2022b), there remains little work doing so for LP. Specifically, we observe that (1) No LP Benchmark Datasets: Current graph benchmark datasets designed with a quantifiable distribution shift are focused solely on the node and graph tasks (Zhou et al., 2022a; Ding et al., 2021), with no datasets available for LP. (2) Absence of Foundational Work: There is limited existing work for distribution shifts relevant to LP (Zhang et al., 2022; Dong et al., 2022). Current methods are primarily focused on detecting and alleviating anomalies within node- and graph-level tasks (Jin et al., 2022; Bevilacqua et al., 2021; Gao et al., 2023; Wu et al., 2024; Guo et al., 2023; Wu et al., 2023; Sui et al., 2024; Li et al., 2022b). Additionally, few methods are designed for aiding LP generalization in any setting (Dong et al., 2022; Zhao et al., 2022; Zhang et al., 2022; Zhou et al., 2022a). Also, other LP generalization methods which are theorized to improve performance in shifted scenarios remain crucially untested (Singh et al., 2021; Wang et al., 2023b).
@@
To tackle these problems, this work proposes the following contributions: (1) Creating Datasets with Meaningful Distribution Shifts. LP requires pairwise structural considerations (Liben-Nowell and Kleinberg, 2003; Mao et al., 2024a). Additionally, when considering realistic settings (Li et al., 2024) or distribution shift (Zhu et al., 2024), GNN4LP models are sensitive to structual changes within datasets relative to graph (Wei et al., 2021; Yuan et al., 2021) and node classification (Shi et al., 2023; Zhao et al., 2023) models. To quantify this sensitivity and better understand distribution shifts, we use key structural LP heuristics to split the links into train/validation/test splits via LPShift. By applying LPShift to generate dataset splits, we induce shifts in the underlying distribution of the links, thereby affecting link formation (Mao et al., 2024a) and downstream performance. This elevates preexisting LP benchmarks by allowing users to control the extent with which their dataset undergoes distribution shift while retaining relevance to LP. Much like how prior benchmarks induce substructure (Gui et al., 2022; Koh et al., 2021) and drug scaffold (Ji et al., 2022) shift to provide relevant challenges within graph and node classification. Theoretical justification is provided in Section 4 and empirical results are provided in Section 5. (2) Benchmarking Current LP Methods. GNN4LP models demonstrate a strong sensitivity to LPShift when generalizing to the synthetic dataset splits. Despite the existence of LP generalization methods, such as FakeEdge (Dong et al., 2022) and Edge Proposal Sets (Singh et al., 2021), there remains little work benchmarking link-prediction models under distribution shifts (Ding et al., 2021; Dong et al., 2022; Zhu et al., 2024). This lack of benchmarking contributes to a gap in understanding, impeding the capabilities of LP models to generalize in real-world scenarios. This work quantifies the performance of SOTA LP models under 16 unique LPShift scenarios and provides analysis as a foundation for improving LP model generalization. We further quantify the effects of LP and graph-specific generalization methods, finding that they also struggle to generalize with differing structural shifts.
@@
The remainder of this paper is structured as follows. In Section 2, we provide background on the heuristics, models, and generalization methods used in LP. In Sections 3 and 4, we detail how structural heuristics, theoretical perspectives, and HeaRT (Li et al., 2024) relate to our proposed splitting strategy. After which, we formally introduce LPShift. Lastly, in Section 5, we benchmark a selection of LP models and generalization methods on LPShift, followed by analysis to understand the effects of this new strategy.
@@ 3. Benchmark Dataset Construction
(a) CN
(b) PA
(c) SP
[…]
3.1. Types of Distribution Shifts
We induce distribution shifts by splitting the links based on key structural properties affecting link formation and thereby LP. We consider three type of metrics: Local structural information, Global structural information, and Preferential Attachment. Recent work by (Mao et al., 2024b) has shown the importance of local and global structural information for LP. Furthermore, due to the scale-free nature of many real-world graphs and how it relates to link formation (Barabási and Albert, 1999), we also consider Preferential Attachment. A representative metric is then chosen for each of the three types, shown as follows:
3.2. Dataset Splitting Strategy
In the last subsection we described the different types of metrics to induce distribution shifts for LP. The metrics cover fundamental structural properties that influence the formation of new links. Given the functionality of: CN, SP, and PA within link prediction (Liben-Nowell and Kleinberg, 2003), LPShift applies the three heuristics to induce distribution shifts which respectively target: local, global, and scale-free properties of the graph. Therefore, providing a meaningfully-difficult challenge to GNN4LP, as empirically-demonstrated in Section 5. We now describe how LPShift applies these heuristic measures to split the dataset into train/validation/test splits to induce such shifts.
[…]
In order to build datasets with structural shift, we apply a given neighborhood heuristic to score each link. This score is then compared to a threshold ($i_{train},i_{valid})$ to categorize a link within a different sample distribution. As denoted in Alg. 1, the heuristic score of the link $(u,v)$ is $h(u,v)$. The link falls into: training when $h(u,v)<i_{train}$, validation when $i_{train}<h(u,v)\leq i_{valid}$, and testing when $h(u,v)>i_{valid}$. The new training graph is constructed from the original OGB dataset (Hu et al., 2020). Validation and testing samples are removed from the new training graph to prevent test-leakage and limited to 100-thousand edges maximum.
@@ 3.2. Dataset Splitting Strategy
7:     $\Psi(u,v)=h(u,v)$triangleright Score edge with heuristic
8:     if $h(u,v)\leq i_{train}$ then
@@
10:     else if $h(u,v)>i_{train}\text{ and }h(u,v)\leq i_{valid}$ then
@@
With Figure 1, we provide a small example of how splits are produced by our proposed splitting strategy. Specifically, Figure 1(a) demonstrates an outcome of the CN split labelled “CN - 1,2” where sampled edges pulled from the: black-dotted line = training (no CNs), red-dotted line = validation, (1 CN), and blue-dotted line = testing ($\geq$2 CNs). Figure 1(b) follows a similar high-level process, focusing on node-degree instead of CNs. As such, the validation (red) cut splits nodes with degree 3 and the testing (blue) cut splits nodes with degree 1. Figure 1(c) details Shortest-Path (SP), where the red cut works on a path of length 3 and is categorized by taking SP as the denominator of the ratio, $\frac{1}{SP}$.
To test how different LPShift thresholds impact performance, we adjust the $i_{train}$ and $i_{valid}$ thresholds to produce 3 varied CN and PA splits; as well as 2 varied SP splits. The variations in split thresholds were chosen based on two conditions. 1) Structural information within splits varies due to user-defined thresholds. In the ”Forward” scenario (i.e. CN - 1,2), splits are given increasingly more structural information between training to validation, giving the model an easier time generalizing under testing. For the ”Backward” scenario (i.e, CN - 2,1), stricter thresholds mean less structural information between validation to testing; making generalization difficult. 2) The final dataset split contains a sufficient number of samples. Each LPShift split requires enough split samples to allow model generalization. Given the limited number of SP split samples, the SP splits were limited to 2 variants.
@@ 3.3. Why use HeaRT with LPShift?
It is common practice for most link prediction baselines to randomly-sample all negatives samples. However, random sampling is shown to be unrealistic, given that it is trivially-easy for a GNN4LP model trained via supervised learning to distinguish positive and negative samples based solely on the sample’s structure (Li et al., 2024). We demonstrate an example of this effect in the right subplot of Figure 2, where the original setting samples negative at random to overwhelmingly contain 0 Common Neighbors. The samples produced by the original setting can be especially misleading, GNN4LP models could achieve high performance due to this overwhelming density of negative samples with 0 CNs versus positive samples with $>$0 CNs. Figure 2’s left subplot shows how HeaRT corrects this problem by sampling valid and test negatives whose distribution of CNs overlap with their respective positive samples (shown in red and blue), as well as the distribution of positive training samples (shown in black on both subplots). Therefore, HeaRT sampling avoids a critical pitfall within the original setting, where GNN4LP models could learn the trivial dynamic where any sample with more than 0 CNs as a positive edge and any sample with 0 CNs as a negative edge.
@@ 4. Why LPShift? A Theoretical Perspective
Section 3 details how heuristic choice and HeaRT influence LPShift’s ability to alter dataset structure. We now consider how GNNs, as neural backbones for GNN4LP, rely on assumptions about the dataset’s underlying sample distributions to learn representations for effective link prediction. From these assumptions, GNN4LP capture structural heuristics associated with link formation, such as LPFormer estimating Katz-Index (Shomer et al., 2024). LPShift undermines these assumptions through direct manipulation of the dataset sample distribution, which is quantifiable within probability theory. Under which, we treat a link predictor as a function to learn a probability measure over a pairwise embedding. This learned probability measure, coupled with the obtained embedding, determines the ability of GNN4LP to quantify whether link formation occurs or not. We provide a theoretical perspective on this effect below:
@@
Indeed, when the i.i.d assumption across dataset samples holds, supervised training of neural LP models with GNN backbones enables message-passing across training samples. This further allows GNN4LP to observe and subsequently model link formation (Bronstein et al., 2021; Zečević et al., 2021), contributing to high downstream performance. However, Theorem 4.2 conditions link formation dynamics within a given graph dataset, meaning that the necessary i.i.d. subsets for learning and then achieving generalized performance are obscured. In effect, the GNN4LP model gets a limited view on pairwise information contained within the dataset. As such, LPShift allows direct testing of the hypothesis: Can GNN4LP model structural heuristics for effective downstream performance in scenarios with limited structural information? We verify this effect empirically in Section 5.
@@ 5. Experiments
To bridge the gap for GNN4LP generalizing under distribution shifts, this work addresses the following questions: (RQ1) Is the distribution shift induced by LPShift significant? (RQ2) Can SOTA GNN4LPs generalize under our proposed distribution shifts? (RQ3) Can current generalization methods boost the performance of GNN models? (RQ4) What components of the proposed distribution shift are affecting the LP model’s performance?
@@ 5.1. Experimental Setup
Datasets: We consider 16 “Forward” and “Backward” LPShift splits for the following 3 OGB datasets (Hu et al., 2020): ppa, ddi, and collab, for a total of 48 tested splits. The resulting datasets represent tasks in three separate domains and three shifted scenarios, allowing a comprehensive study of LP generalization under distribution shift. For all datasets, we create multiple splits corresponding to each structural property, as detailed in Section 3.2. For the “Forward” split, denoted as $(X,Y,Z)$, an increase in $Y$ and $Z$ indicates more structural information available to the training adjacency matrix. The “Backward” split swaps the training and testing splits from their counterpart in the “Forward” split, resulting in the training adjacency matrix losing access to structural information as $X$ and $Y$ increase.
@@
Graph-Specific Generalization Methods: We also test the performance of BUDDY with multiple generalization techniques. This includes DropEdge (Rong et al., 2020), which randomly removes a portion of edges from the training adjacency matrix. Edge Proposal Sets (EPS) (Singh et al., 2021), which uses paired LP models to filter and rank top-k edges to insert as Common Neighbors in the training adjacency matrix. Lastly, we consider Topological Concentration (TC) (Wang et al., 2023b), which re-weights the edges within the training adjacency matrix based on their shared information. Note: We were unable to test CFLP (Zhao et al., 2022) and EERM (Wu et al., 2022) after adapting their current implementations to the HeaRT evaluation setting. CFLP experienced an out-of-memory error on all tested LPShift splits. EERM experienced an out-of-memory error on every LPShift split of ogbl-ppa and ogbl-ddi, while also exceeding 48 hours per run on ogbl-collab without converging.
Traditional Generalization Methods: The final round of tests applies: IRM (Arjovsky et al., 2019), VREx (Krueger et al., 2021), GroupDRO (Sagawa et al., 2019), DANN (Ganin et al., 2016), and DeepCORAL (Sun and Saenko, 2016) as generalization methods to GCN. Given that LPShift splits graph data directly on the sample structure, all tests with traditional generalization methods treat training samples as an environmental partition. The partitions =$\mathcal{E}=\{e_{1},...,e_{N}\}$ are composed of $N$ environmental subsets, where each subset is binned based on their neighborhood heuristic; as demonstrated for CNs in Figure 3. This binning of environmental subsets forces traditional generalization methods to target graph structure, controlling for the method’s ability to improve GNN4LP performance under structural shift induced by LPShift.
Figure 3. The original environment of dataset samples (shown left) split into subsets by their CN heuristic scores (shown right).
@@
The original environment of dataset samples (shown left) split into subsets by their CN heuristic scores (shown right).
@@
Hyperparameters: All methods were tuned on permutations of learning rates in $\{1e^{-2},1e^{-3}\}$ and dropout in $\{0.1,0.3\}$. Each model was trained and tested over five seeds to obtain the mean and standard deviations of their results. Given the significant time complexity of training and testing on LPShift variants for the ogbl-ppa datasets, NCNC and LPFormer were tuned on a single seed, followed by evaluation of the tuned model on five separate seeds. Additional hyperparameter tuning details are included in Appendix A.
5.2. Results for GNN4LP
In order to provide a unified perspective on how distribution shift affects link prediction models, GNN4LP models were trained and tested across five seeded runs on versions of ogbl-collab, ppa, and ddi split by: Common Neighbors, Shortest-Path, and Preferential-Attachment. Examining the results, we have the following three key observations.
Observation 1: Drastic Changes in Model Performance. As shown in Table 1, the majority of synthetic splits generated by LPShift reduce model performance significantly below the original HeaRT standard (Li et al., 2024). This is especially pronounced in denser graphs like ogbl-ppa and ogbl-ddi, indicating that LPShift affects the varied structures of graph data, which then reduces downstream perfomance. The only exception is ogbl-collab, where: RA, NCNC, LPFormer, and SEAL experience measurable improvements over the HeaRT baseline. Given ogbl-collab’s lack of density relative to ogbl-ppa and ogbl-ddi, there are considerations about how LPShift induces a covariate shift to features (Koh et al., 2021), especially where structure is correlated to the feature distribution.
Additional analysis shown in Figure 6 indicates how dataset structure relates to the performance effects of LPShift.
Observation 2: Performance Differs Across LPShift Splits. As shown in Figure 4, regardless of whether a model is tested on a “Forward” or “Backward” split of ogbl-ppa; the change in structural information for each subsequent split gradually changes a model’s performance. On the “CN” split, a stark increase is seen between ”Split 1” and ”Split 2” (also detailed in descending order within Table 10), indicating that more structural information between training and validation improves GNN4LP performance (Wang et al., 2023a), even on dense datasets like ogbl-ppa. The fact that these results include splits produced via Preferential-Attachment, Global Structural Information (SP), and Local Structural Information (CN) indicates the effect of any change in structural information when training LP models (Mao et al., 2024b). In principle, Observation 1 and 2 answer RQ1 and RQ2 by indicating that structural alignment between positive samples from separate splits is critical for GNN4LP performance. LPShift makes this effect quantifiable through user-controlled thresholds, enabling future research to: induce, diagnose, and correct problems caused by LP-specific distribution shift.